# 6G comprehensive intelligence: network operations and optimization based on Large Language Models

### v1

@@ V Several open issues
With the continuous development and improvement of 6G networks, the application of Large Language Models within them will become more extensive and profound. However, we also face numerous challenges. On the one hand, as the scale and complexity of models increase, the costs of training and inference will also rise. On the other hand, it is an urgent issue to achieve efficient model operation and performance optimization while ensuring data privacy and security. Therefore, we need to continuously explore new technologies and methods to address these challenges. In summary, the efficient network operation and performance optimization of LLM in 6G networks are a research area with significant importance. Through in-depth research and exploration, we hope to provide more intelligent and efficient LLM support for the development and application of 6G networks, thus driving the process of comprehensive intelligence in human society. However, important issues will also be faced in the research of intelligent networks and performance optimization.
V-A Scaling problem
The model has high requirements for computing and storage resources, making it difficult to deploy collaboratively at the edge. How to set up a Cloud-Edge-Device collaborative framework to properly allocate resources and coordinate capabilities has become a new challenge. In addition, model lightweighting has become a solution to successfully deploying at the edge, and exploring lightweight models has become another important research topic.
V-B Multimodal alignment problem
@@
V-C Selection and equilibrium of LLM and small model
@@
V-D Model self evolution problem
@@
V-E Data collection and real-time performance of the system
Large-scale data collection faces a series of issues, including the costs of data transmission and storage, as well as concerns regarding data accuracy, reliability, privacy, and security protection. At the same time, data security issues should not be ignored. For example, Mišić et al. [15] discussed in detail a large number of existing security and privacy related restrictions on the network, which reflected the importance of network security in the context of Internet of things applications. To address these challenges, it is necessary to adopt advanced data compression and encryption technologies while leveraging cloud computing and edge computing techniques to achieve efficient data processing and utilization. However, due to the complexity and uncertainty of network transmission, as well as the limitations of device processing capabilities, achieving system real-time performance is not an easy task. Therefore, one reliable approach is to incorporate data simulation and future prediction from digital twin networks to tackle this difficulty.
V-F Security and interpretability of LLM in networks
@@ VII Acknowledgments
This work was supported by the Postdoctoral Fellowship Program of CPSF (Grant No.GZC20233160), Changsha Municipal Natural Science Foundation (Grant no.kq2208284), Hunan Provincial Natural Science Foundation (Grant no.2023jj40774), National Natural Science Foundation of China (Grant no.62302527), and the High Performance Computing Center of Central South University.

### v3

@@ IV Our demo
It is noteworthy that in the architecture diagram depicted in Figure 5, drones communicate information via a dedicated network, ensuring efficient real-time data transmission. Furthermore, time-series data incorporating network information is first processed through a large language model (LLM) for semantic understanding and conversion into textual information. Simultaneously, logical rules are employed to symbolize signal data, creating fused data from multiple modalities for subsequent LLM training. This data encompasses flight control information and various related parameters of drones, potentially containing fault samples as well. For instance, a circuit malfunction in a drone may trigger a series of chain reactions leading to data anomalies.
Moreover, a knowledge graph is utilized to extract and unify data from these different modalities, which is then fed into the LLM for training. Fine-tuning of the LLM is also conducted as necessary. Finally, the large model conducts health assessments based on the input data. It trains a fault discriminator through embedding models and can also leverage causal inference methods in conjunction with the LLM for troubleshooting. Subsequently, a health report is generated, outlining fault types, faulty nodes, and corresponding solutions, thereby achieving comprehensive network health assessment for the entire system.
In summary, the design and implementation of a drone health assessment system based on Large Language Models necessitate comprehensive consideration of multiple aspects, including data quality and bias, model training and validation, security and privacy protection, system architecture and integration, real-time performance and accuracy, user-friendliness, as well as compliance and ethical issues. By continually optimizing and refining these aspects, it can enhance the performance and reliability of the drone health assessment system, thereby providing robust support for the safe flight of drones.
@@ V Several open issues
With the continuous development and improvement of 6G networks, the application of Large Language Models within them will become more extensive and profound. However, we also face numerous challenges. On the one hand, as the scale and complexity of models increase, the costs of training and inference will also rise. On the other hand, it is an urgent issue to achieve efficient model operation and performance optimization while ensuring data privacy and security. Therefore, we need to continuously explore new technologies and methods to address these challenges. However, important issues will also be faced in the research of intelligent networks and performance optimization.
V-A Model generalization ability and robustness
The network performance optimization system based on Large Language Models may have the problem of insufficient generalization ability. Although Large Language Models have strong learning capabilities, they may suffer from overfitting and insufficient generalization ability in some cases. This means that the model may not adapt well to new operation scenarios and tasks, limiting its effectiveness in practical applications. Moreover, the issue of robustness is also a key technology that needs to be urgently addressed. The network health assessment system based on Large Language Models needs to handle various exceptional situations, including unknown faults, system crashes, and so on. Large language models need to possess sufficient robustness to cope with these challenges.
V-B Data understanding and representation
In the context of network operation and performance optimization leveraging LLMs, understanding the representation of multimodal data such as semantics, text, images, and videos is a complex yet crucial process that significantly contributes to providing decision support for network operation. However, there remain significant challenges in exploring and implementing universal methods for the effective fusion and comprehension of multimodal data through technical means such as data alignment and correlation, cross-modal feature representation, and semantic understanding and reasoning.
Different modalities of network data (e.g., text, images, videos) may exhibit temporal or spatial correspondences. To effectively fuse these data, algorithms with superior performance need to be designed for data alignment and correlation operations, ensuring that data from various modalities can complement and enhance each other.
Furthermore, to uniformly represent the diverse modalities of network data within the same feature space, innovative cross-modal feature representation techniques are required. These techniques convert data from different modalities such as text, images, and videos into shared feature vectors or representations, enabling further processing and analysis by LLMs.
Once a unified representation of multimodal data is achieved, a more powerful LLM framework with enhanced representation capabilities must be designed for semantic understanding and reasoning. This framework should be capable of comprehending semantic information in network log texts, identifying key elements and events in images and videos, and inferring associations and causal relationships among different modalities of network operation data.
V-C Multimodal alignment problem
@@
V-D Selection and equilibrium of LLM and small model
@@
V-E Model self evolution problem
@@
V-F Data collection and real-time performance of the system
Large-scale data collection faces a series of issues, including the costs of data transmission and storage, as well as concerns regarding data accuracy, reliability, privacy, and security protection. At the same time, data security issues should not be ignored. For example, Mišić et al. [15] discussed in detail a large number of existing security and privacy related restrictions on the network, which reflected the importance of network security in the context of Internet of things applications. To address these challenges, it is necessary to adopt advanced data compression and encryption technologies while leveraging cloud computing and edge computing techniques to achieve efficient data processing and utilization, this can effectively utilize the latest computing modes to solve the problems caused by data transmission. However, due to the complexity and uncertainty of network transmission, as well as the limitations of device processing capabilities, achieving system real-time performance is not an easy task. Therefore, one reliable approach is to incorporate data simulation and future prediction from digital twin networks to tackle this difficulty.
V-G Security and interpretability of LLM in networks
@@ VII Acknowledgments
This work was supported by the Postdoctoral Fellowship Program of CPSF (Grant No.GZC20233160), Changsha Municipal Natural Science Foundation (Grant no.kq2208284), Hunan Provincial Natural Science Foundation (Grant no.2023jj40774), National Natural Science Foundation of China (Grant no.62302527), Hunan Provincial Natural Science Foundation Enterprise Joint Fund (Grant no.2024JJ9173). We sincerely appreciate the efforts made by the editor and reviewers to improve this paper.